OpenStudy (anonymous):

4. Divide: b. 13x = 39

OpenStudy (saifoo.khan):

x = 39/13 x = 3

myininaya (myininaya):

divide both sides by 13 13/13=1 so 13x/13=1*x=x so we have x=39/13, so x=3

OpenStudy (saifoo.khan):

Thnks! :)

myininaya (myininaya):

:)

OpenStudy (saifoo.khan):

[:

myininaya (myininaya):

master khan!

OpenStudy (saifoo.khan):

hehe Thnks! :)

OpenStudy (saifoo.khan):

myininaya, whts your real name?

OpenStudy (anonymous):

SO DO YOU JUST DIVIDE THE SMALLER NUMBER ON BOTH SIDES?

OpenStudy (saifoo.khan):

i take a shortcut, cross multiply them. ^.^

OpenStudy (anonymous):

BUT CAN YOU DO IT THAT WAY?

OpenStudy (saifoo.khan):

Yeah, u can.. but the best way is the way u knw.

myininaya (myininaya):

christy

OpenStudy (saifoo.khan):

Christy?

myininaya (myininaya):

if you have ax=b and you are solving for x you always divide both sides by a

myininaya (myininaya):

christy=my name

OpenStudy (saifoo.khan):

Lol, okay.. i got that.. :)